# ParaCook: On Time-Efficient Planning for Multi-Agent Systems

### v1

@@ 1 Introduction
Large Language Models (LLMs) have empowered agents with remarkable planning capabilities in complex and interactive tasks Yang et al. (2025); Shinn et al. (2023). Planning integrates task knowledge and breaks down long-horizon goals into subtasks, enabling overall coherence rather than just local optimality. However, time efficiency remains a major challenge, as it is widely accepted that the usability of LLM agents depends not only on improvements in raw model intelligence or accuracy, but on maximizing Agentic ROI Liu et al. (2025).
@@
This investigates the research question that long-horizon tasks should not only be distributed correctly, but also scheduled efficiently. Efficient scheduling is non-trivial because complex tasks always involve both parallel and sequential steps. For a single agent, its schedule can exploit concurrency between sub-tasks that do not block each other to reduce idle time, such as during water boiling, the agent can switch to another action like chopping vegetables. For multi-agent system, different agents can leverage the collaboration to distribute workload to operate on independent subtasks, such as cooking different dishes. These two forms of parallelism, intra- and inter-agent, jointly determine the time efficiency of multi-agent planning.
@@
However, existing agent benchmarks mainly focus on correctness and rule-following rather than time efficiency in parallel scheduling. While several studies have considered concurrent scenarios, their coverage remains limited. AsyncHow (Lin et al., 2024a) oversimplifies tasks and assumes unlimited resources. Robotouille (Gonzalez-Pumariega et al., 2025) focuses only on single-agent asynchronous planning. CookBench (Cai et al., 2025) offers long-horizon, multi-step tasks, but is too comprehensive to isolate time efficiency evaluation. As a result, current benchmarks do not fully expose opportunities for parallelism, leaving a notable gap in evaluating whether LLMs can truly exploit concurrency to minimize overall task completion time.
@@
To address this gap, we propose ParaCook, a new benchmark that focuses on evaluating whether LLMs can effectively schedule tasks and coordinate agents to minimize overall completion time. We focus on cooking, an everyday scenario that naturally involves both sequential and parallel tasks.Inspired by the Overcooked game, ParaCook simplifies action spaces and task rules, allowing models to concentrate on effective task scheduling and optimal time utilization. Our benchmark supports multiple agents cooperating, emphasizing both inter-agent parallelism and intra-agent time utilization. To further isolate planning from low-level actions, we introduce corresponding abstract tasks that focus solely on high-level task allocation and optimization of execution time. These abstract tasks enable direct evaluation of LLMs’ planning capabilities and preserve the challenges of efficient multi-agent scheduling.
@@ 2 Related Work
In this section, we introduce relevant studies on LLM as planning agents, their extension to multi-agent cooperation, and benchmarks for evaluating planning quality.
@@ 2.1 Planning for Agents
LLMs have emerged as powerful planner due to their strong reasoning capabilities.
@@
Multi-agent system planning. Multi-agent systems (MAS) have shown great progress in complex tasks, where agents with diverse profiles collaborate. Planning becomes even more pronounced in MAS, requiring collaboration and ROI (Ma et al., 2025). A commander, meta-agent, or manager agent is often responsible for iter-agent planning, like role assignment, and coordination among agents. Early frameworks (Wu et al., 2023; Hong et al., 2024) established structured workflows under centralized control. More flexible planning strategies encourage hierarchical and partially decentralized planning, as in AgentOrchestra (Zhang et al., 2025c) and Agent-Oriented Planning (Li et al., 2025).
@@ 2.2 Parallelable Agent
Recent efforts have explored making LLM-based agents parallelizable to enhance efficiency. Adaptive reasoning frameworks enable concurrent thought processes (Pan et al., 2025; Zhang et al., 2024), while DAG-based and divide–aggregate methods support simultaneous tool use (Zhang et al., 2025a; Zhu et al., 2025). At the system level, asynchronous planning–acting architectures and graph-based schedulers (Zhang et al., 2025b) facilitate concurrent execution and coordination. Despite these advances, research remains disparate across reasoning and scheduling, with no unified benchmark for evaluating agents’ ability to exploit parallelism effectively.
@@ 2.3 Benchmarks for Agent Planning
Rigorous benchmarks are essential for systematic evaluation, especially in multi-agent settings. Most existing benchmarks focus on single-agent or strictly sequential planning, emphasizing solution correctness and rule-following. AsyncHow (Lin et al., 2024b) evaluates asynchronous planning by computing an optimal schedule, but assumes unlimited resources to achieve maximal parallelism, and lacks environment interaction. Robotouille (Gonzalez-Pumariega et al., 2025) simulates cooking tasks with proper time consumption, but only focuses on single-agent asynchronous planning. WORFBENCH (Qiao et al., 2024) decomposes tasks into workflow graphs but evaluates graph similarity rather than execution time. Collab-Overcooked (Sun et al., 2025)and CookBench (Cai et al., 2025) provide comprehensive challenges for MAS. However, they also do not support explicit time efficiency evaluation. As is summarized in Table 1, existing benchmarks still fall short of systematically testing how effectively LLMs can exploit both intra-agent and inter-agent parallelism to minimize overall completion time, motivating our design of ParaCook.
@@ 4.1 Environment
The environment of ParaCook is a 2D grid world that simulates a kitchen, supporting agent–environment and agent–agent interactions while providing feedback. At the beginning of each episode, the environment initializes an agent team, the task to be performed, and the map configuration.
@@
  • Process. Perform a continuous operation at a workstation, such as chopping or washing. The execution time depends on the specific operation.
@@ 5 Metrics
ParaCook evaluates plans along two primary dimensions: correctness and efficiency. Correctness is measured by the dataset-level Success Rate (SR), while efficiency is primarily captured by the task-level Order Completion Time (OCT). Additional metrics, including Movement Distance (MD) and Agent Utilization (AU), provide fine-grained analyses of execution cost and coordination.
@@ 5.1 Primary Metrics
where $T_{i}^{\ast}=OCT_{i}$ if task $i$ succeeds, and $T_{i}^{\ast}=T_{\max}^{i}$ otherwise. Here, $T_{\max}^{i}$ is the predefined upper bound for task $i$. pOCTpOCT penalizes failures with maximal time, while $nOCT$ normalizes efficiency among successful runs.
@@ 5.2 Auxiliary Metrics
with $D_{i}^{\ast}=MD_{i}$ for successful tasks and $D_{i}^{\ast}=D_{\max}^{i}$ otherwise. $D_{\max}^{i}$ represents the predefined upper bound of movement distance for task $i$.
@@ 6.1 Experimental Setup
We manually annotated six categories of recipes with increasing levels of difficulty, and each category contains several different dishes. For orders, a set of dishes was randomly sampled from a given recipe category, with the number of dishes ranging from one to four to control the complexity of the task. To encourage parallelism, each kitchen map was equipped with two stations of each type, placed randomly while ensuring connectivity. The number of agents varied between one and three. For each configuration, we used five random seeds to generate different orders and maps.
@@
To evaluate different methods, we consider the I/O and CoT prompting methods. I/O produces a complete plan directly from the initial state, which specifies the actions for each agent. CoT (Wei et al., 2022) encourages reasoning while planning, explicitly predicting intermediate states. We evaluate state-of-the-art API LLMs, including GPT-5, Claude-Opus-4.1, Gemini-2.5-Pro Comanici et al. (2025), DeepSeek-V3.2-Exp DeepSeek-AI (2025), and Qwen3-Max-Preview.
Success Rate (SR, %) Penalized Order Completion Time (pOCT)
Model Easy Medium Hard Avg. Easy Medium Hard Avg.
GPT-5 80.83 69.17 45.00 65.00 137.98 330.80 416.35 295.04
Gemini-2.5-Pro 60.00 55.07 27.14 47.40-17.60 191.94 390.33 486.94 356.40+61.36
DeepSeek-V3.2-Exp 66.67 47.50 21.67 45.28-19.72 176.89 467.98 521.19 388.69+93.65
Claude-Opus-4.1 26.67 12.50 0.00 13.06-51.94 268.54 585.65
Qwen3-Max-Preview 6.67 0.00 0.00 2.22-62.78 291.97
Table 2: Success Rate (SR) and penalized Order Completion Time (pOCT) of different models across three difficulty levels and their averages. Bold indicates the best performance in each column. Red subscripts in the Avg. column show the performance gap relative to the best model (GPT-5).
@@ 6.2 Results
In terms of pOCT, GPT-5 again demonstrates the best performance, with the lowest average planning overhead (295.0), consistently faster than other models at all difficulty levels. Gemini-2.5-Pro (356.4) and DeepSeek-V3.2-Exp (388.7) occupy the middle range, showing reasonable efficiency but still substantially slower than GPT-5. Claude-Opus-4.1 (356.7) and Qwen3 (289.4) exhibit unstable efficiency: despite occasionally achieving low overhead on Easy tasks, they either fail to complete tasks at higher difficulty levels or produce unreliable results, limiting their practical applicability.
In detail, Table 5 illustrates that time efficiency is related to nOCT, MD and AU. Specifically, GPT-5’s superior efficiency is reflected in strong auxiliary metrics, maintaining minimal nOCT and pMD across all tasks, and high agent utilization (AU) above 86%. Gemini-2.5-Pro and DeepSeek-V3.2-Exp are less efficient, with pMD values rising to 178.60 and 195.28 on Medium tasks, respectively. Claude-Opus-4.1 performs the worst, with pMD exceeding 230 and significantly lower utilization, confirming that inefficient scheduling and the failure to parallelize actions directly lead to longer completion times.
@@
Figure 2: Results on the human-evaluated subset for LLM-human comparison. Detailed scores are in Table 6 and 7.
Figure 2 compares human participants with top-performing LLMs on a curated subset of tasks (whose detailed results are presented in Table 6). Humans achieve a perfect 100% success rate across all difficulty levels, highlighting their robustness in complex, long-horizon scenarios where LLMs may fail. The performance gap becomes particularly evident on Hard tasks. For instance, the top-performing model, GPT-5 with CoT, achieved a 57% SR, whereas humans made no errors.
In terms of efficiency, while top LLMs are competitive on easy tasks, they lag significantly as complexity increases. On Hard tasks, humans completed orders with an nOCT of 15.38, whereas GPT-5 required a longer time of 17.61. The disparity in spatial efficiency is even more stark, with humans having a pMD of just 47.75, compared to GPT-5’s 143.55, indicating far more optimized movement from humans. Interestingly, AU is slightly lower for humans (87.13%) than for GPT-5 (94.34%). This is not due to inefficiency, but because human plans are executed so quickly that the necessary waiting time between dependent tasks constitutes a larger proportion of the total duration. Overall, these results demonstrate that while LLMs can produce correct and parallelized plans for less complex tasks, their execution strategies remain less temporally and spatially optimized than those of humans, particularly for high coordination demands.
@@ 7.1 Performance across Task Complexity
Next, we examine the effect of the planning horizon. As anticipated, increasing the number of orders extends the task sequence, leading to a general decline in SR (Figure (b1)). Consequently, the overall cost metrics, pOCT (Figure (b2)) and pMD (Figure (b4)), rise steadily as more work requires more time and movement. The nOCT trend is more nuanced (Figure (b3)); the metric remains stable or even decreases when moving from one to two orders. This is because sufficient initial resources allow agents to prepare two orders in parallel, so the actual completion time $T_{\text{actual}}$ increases only modestly. In contrast, the normalization factor, $T_{\max}$, is set significantly higher for a two-order task. This combination results in a stable or smaller nOCT ratio, indicating scalable plan quality in scenarios with low complexity. Lastly, AU generally improves with more orders (Figure (b5)), as more subtasks effectively reduce agent idle time in successful runs.
@@
GPT-5 IO 80.83 137.98 69.17 330.80 45.00 416.35 65.00 295.04
CoT 84.17+3.34 129.59-8.39 77.03+7.86 283.03-47.77 57.39+12.39 359.16-57.19 72.86+7.86 257.26-37.78
Gemini-2.5-Pro IO 60.00 191.94 55.07 390.33 27.14 486.94 47.40 356.40
CoT 55.83-4.17 201.43+9.49 47.22-7.85 419.10+28.77 37.50+10.36 447.49-39.45 46.85-0.55 356.01-0.39
DeepSeek-V3.2-Exp IO 66.67 176.89 47.50 467.98 21.67 521.19 45.28 388.69
CoT 67.50+0.83 178.42+1.53 45.83-1.67 470.74+2.76 40.83+19.16 465.84-55.35 51.39+6.11 371.67-17.02
Claude-Opus-4.1 IO 26.67 268.54 12.50 585.65 0.00 13.06 427.10
CoT 55.00+28.33 215.99-52.55 14.58+2.08 577.52-8.13 0.00+0.00 23.19+10.13 396.76-30.34
Qwen3-Max-Preview IO 6.67 291.97 0.00 0.00 2.22 291.97
CoT 10.83+4.16 286.73-5.24 0.00+0.00 0.00+0.00 3.61+1.39 286.73-5.24
Table 3: Results with CoT. Blue subscripts indicate improvements (higher SR or lower pOCT), red indicates degradation. The $\delta$(CoT-IO) row shows the difference between CoT and IO methods.
7.2 Performance with CoT

### v2

@@ 1 Introduction
Large Language Models (LLMs) have empowered agents with remarkable planning capabilities in complex and interactive tasks Yang et al. (2025); Shinn et al. (2023). Planning integrates task knowledge and breaks down long-horizon goals into subtasks, enabling overall coherence rather than just local optimality. However, in long-horizon tasks with explicit temporal dynamics and concurrency, task correctness alone is no longer sufficient to characterize planning quality. Even when all required subtasks are completed successfully, different plans may exhibit vastly different execution times due to idle periods, blocking, and resource contention. As a result, effective planning in such settings fundamentally requires reasoning about time-efficient scheduling, rather than correctness alone.
@@
This motivates the research question that long-horizon tasks should not only be distributed correctly, but also scheduled efficiently. Efficient scheduling is non-trivial because complex tasks inherently involve both parallel and sequential steps. For a single agent, schedules can exploit intra-agent parallelism between sub-tasks that do not block each other to reduce idle time. For example, during water boiling, the agent can switch to another action, like chopping vegetables. For multi-agent system, different agents can further leverage inter-agent parallelism by distributing independent subtasks, such as cooking different dishes simultaneously. Crucially, these two forms of parallelism jointly determine planning quality under parallel execution, where multiple success-equivalent plans may differ significantly in overall completion time.
@@
From this perspective, existing agent benchmarks are primarily designed to evaluate objectives orthogonal to time-efficient parallel scheduling. Many focus on task success, rule-following, or coordination feasibility, without explicitly modeling execution time or shared-resource contention. Other benchmarks consider concurrency under highly simplified assumptions. AsyncHow (Lin et al., 2024a) reduces planning to ordering pre-decomposed subtasks with given durations and dependency graphs, rather than execution-grounded scheduling. Robotouille (Gonzalez-Pumariega et al., 2025) studies asynchronous execution but is limited to single-agent planning. CookBench (Cai et al., 2025) provides long-horizon embodied tasks, but its comprehensive visual and interaction complexity makes it difficult to isolate and evaluate scheduling efficiency. As a result, current benchmarks do not fully expose opportunities for parallelism, leaving a notable gap in evaluating whether LLMs can truly exploit concurrency to minimize overall task completion time.
@@
To address this gap, we propose ParaCook, a benchmark designed to evaluate time-efficient parallel scheduling in embodied multi-agent planning. We focus on cooking, an everyday scenario that naturally involves both sequential and parallel tasks. Inspired by the Overcooked game, ParaCook simplifies action spaces and task rules, allowing models to concentrate on effective task scheduling and optimal time utilization. Our benchmark supports multiple agents cooperating, emphasizing both inter-agent parallelism and intra-agent time utilization. To further isolate planning from low-level actions, we introduce corresponding abstract tasks that focus solely on high-level task allocation and optimization of execution time, enabling direct evaluation of LLMs’ planning capabilities and preserve the challenges of efficient scheduling.
@@ 2 Related Work
In this section, we introduce studies on agents’ planning, their extension to multi-agent cooperation, and benchmarks for evaluating planning.
@@ 2.1 Planning for Agents
LLMs have emerged as powerful planners due to their strong reasoning capabilities.
@@
Multi-agent system planning. Multi-agent systems (MAS) exhibit great progress in complex tasks, where agents with diverse profiles collaborate, where planning becomes even more pronounced. Early frameworks (Wu et al., 2023; Hong et al., 2024) established structured workflows under centralized control, where a meta-agent is responsible for inter-agent planning, like role assignment and coordination. Recently, flexible planning strategies encourage hierarchical and partially decentralized planning Zhang et al. (2025c); Li et al. (2025).
@@ 2.2 Parallelable Agent
Recent efforts have explored making LLM-based agents parallelizable to enhance efficiency. Adaptive reasoning frameworks enable concurrent thought processes (Pan et al., 2025; Zhang et al., 2024a), while DAG-based and divide–aggregate methods support simultaneous tool use (Zhang et al., 2025a; Zhu et al., 2025). At the system level, asynchronous planning–acting architectures and graph-based schedulers (Zhang et al., 2025b) facilitate concurrent execution and coordination. Despite these advances, research remains disparate across reasoning and scheduling, with no unified benchmark for evaluating agents’ ability to exploit parallelism effectively.
@@ 2.3 Benchmarks for Agent Planning
TimeArena $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ Short
@@
Overcooked-AI $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ Long
@@
Rigorous benchmarks are essential for systematic evaluation, especially in multi-agent settings. However, most existing benchmarks are not designed to evaluate time-efficient parallel scheduling under execution time constraints, and therefore cannot assess how effectively agents exploit parallelism to minimize overall completion time. As summarized in Table 1, none of these benchmarks jointly evaluate intra- and inter-agent parallelism, and time efficiency under realistic execution dynamics. ParaCook is designed to fill this gap by explicitly modeling execution time, asynchronous waiting, and shared-resource contention in a multi-agent embodied environment. Detailed comparisons with benchmarks are provided in Appendix A.
@@ 4.1 Environment
The environment of ParaCook is a 2D grid world that simulates a kitchen, supporting agent –environment and agent–agent interactions while providing feedback. At the beginning of an episode, the environment initializes an agent team, the task to be performed, and the map configuration.
@@
  • Process. Perform a continuous operation at a workstation, such as chopping. The execution time depends on the specific operation.
@@
To isolate parallel scheduling from path-finding complexity, agents can overlap during movement without collisions, and item transfers occur via interaction with surfaces rather than direct handoffs. Action execution times are fixed and summarized in Appendix D.
@@ 5 Metrics
ParaCook evaluates plans primarily along correctness and efficiency. This section details Success Rate for correctness, Order Completion Time for efficiency, and auxiliary metrics including Movement Distance and Agent Utilization.
@@ 5.1 Primary Metrics
where $T_{i}^{\ast}=OCT_{i}$ if task $i$ succeeds, and $T_{i}^{\ast}=T_{\max}^{i}$ otherwise. Here, $T_{\max}^{i}$ is the predefined upper bound for task $i$, details are provided in Appendix G. $pOCT$ penalizes failures with maximal time, while $nOCT$ normalizes efficiency among successful runs.
@@ 5.2 Auxiliary Metrics
with $D_{i}^{\ast}=MD_{i}$ for successful tasks and $D_{i}^{\ast}=D_{\max}^{i}$ otherwise. $D_{\max}^{i}$ represents the predefined upper bound of movement distance for task $i$, details are provided in Appendix G.
@@ 6.1 Experimental Setup
We manually annotated six categories of recipes with increasing levels of difficulty, and each category contains several different dishes. For orders, a set of dishes was randomly sampled from a given recipe category, with the number of dishes ranging from one to four to control the complexity of the task. To encourage parallelism, each kitchen map was equipped with two stations of each type, placed randomly while ensuring connectivity. The number of agents varied between one and three. For each configuration, we used five random seeds to generate different orders and maps. All evaluation configurations and instance counts are enumerated in Appendix H.
@@
We evaluate different centralized planning methods, where a single LLM planner generates actions for all agents. We consider the I/O and CoT prompting methods. I/O produces a complete plan directly from the initial state and specifies the actions for each agent. CoT (Wei et al., 2022) encourages reasoning while planning, explicitly predicting intermediate states. We evaluate state-of-the-art API LLMs, including GPT-5, Claude-Opus-4.1, Gemini-2.5-Pro Comanici et al. (2025), DeepSeek-V3.2-Exp DeepSeek-AI (2025), and Qwen3-Max-Preview.
Success Rate (SR, %) Penalized OCT (pOCT) Normalized OCT (nOCT)
Model Easy Medium Hard Avg. Easy Medium Hard Avg. Easy Medium Hard Avg.
[…]
Table 2: Success Rate (SR), Penalized and Normalized Order Completion Time (pOCT, nOCT) of different models across three difficulty levels and averages. Values shown as mean $\pm$ 95% confidence interval. Bold indicates the best performance. Red subscripts in the Avg. column show the performance gap relative to the best model (GPT-5).
@@ 6.2 Results
In terms of pOCT, GPT-5 demonstrates the best performance, with the lowest average planning overhead (295.0), consistently faster than other models at all difficulty levels. Gemini-2.5-Pro (356.4) and DeepSeek-V3.2-Exp (388.7) occupy the middle range, showing reasonable efficiency but still substantially slower than GPT-5. Examining efficiency on successful runs (nOCT), GPT-5 achieves the best average of 27.29, with DeepSeek-V3.2-Exp close behind at 27.59, while Gemini-2.5-Pro lags at 29.83. This indicates that even among successful plans, GPT-5 consistently produces more time-efficient schedules. Claude-Opus-4.1 and Qwen3 exhibit unstable performance: despite occasionally achieving low overhead on Easy tasks, they either fail to complete tasks at higher difficulty levels or produce unreliable results, limiting their practical applicability.
In detail, Table 5 illustrates that time efficiency is also reflected in MD and AU. Specifically, GPT-5 maintains minimal pMD across all tasks and high agent utilization (AU) above 86%. Gemini-2.5-Pro and DeepSeek-V3.2-Exp are less efficient, with pMD values rising to 178.60 and 195.28 on Medium tasks, respectively. Claude-Opus-4.1 performs the worst, with pMD exceeding 230 and significantly lower utilization, confirming that inefficient scheduling and the failure to parallelize actions directly lead to longer completion times.
@@
Figure 2: Human-evaluated subset for LLM-human comparison. Detailed scores are in Table 6 and 7.
Figure 2 compares human participants with top-performing LLMs on a curated subset of tasks (whose detailed results are presented in Table 6). Details of the human evaluation protocol and interface are provided in Appendix M. Humans achieve a perfect 100% success rate across all difficulty levels, highlighting their robustness in complex, long-horizon scenarios where LLMs may fail. The performance gap becomes particularly evident on Hard tasks. For instance, the top-performing model, GPT-5 with CoT, achieved a 57% SR, whereas humans made no errors.
In terms of efficiency, while top LLMs are competitive on easy tasks, they lag significantly as complexity increases. On Hard tasks, humans completed orders with an nOCT of 15.38, whereas GPT-5 required a longer time of 17.61. The disparity in spatial efficiency is even more stark, with humans having a pMD of just 47.75, compared to GPT-5’s 143.55, indicating far more optimized movement from humans. Interestingly, AU is slightly lower for humans (87.13%) than for GPT-5 (94.34%). This is not due to inefficiency, but because human plans are executed so quickly that the necessary waiting time between dependent tasks constitutes a larger proportion of the total duration. Overall, these results demonstrate that while LLMs can produce correct and parallelized plans for less complex tasks, their execution strategies remain less optimized than those of humans, particularly for high coordination demands. A detailed time-budget decomposition is provided in Appendix I.
@@ 7.1 Performance across Task Complexity
In terms of the planning horizon, increasing the number of orders extends the task sequence, leading to a general decline in SR (Figure (b1)). Consequently, the overall cost metrics, pOCT (Figure (b2)) and pMD (Figure (b4)), rise steadily as more work requires more time and movement. The nOCT trend is more nuanced (Figure (b3)); the metric remains stable or even decreases when moving from one to two orders. This is because sufficient initial resources allow agents to prepare two orders in parallel, so the actual completion time $T_{\text{actual}}$ increases only modestly. In contrast, the normalization factor, $T_{\max}$, is set significantly higher for a two-order task. This combination results in a stable or smaller nOCT ratio, indicating scalable plan quality in scenarios with low complexity. Lastly, AU generally improves with more orders (Figure (b5)), as more subtasks effectively reduce agent idle time in successful runs.
@@
[…]
Table 3: Results with CoT. Values are reported as mean $\pm$ 95% confidence interval. Blue subscripts indicate improvements (higher SR or lower pOCT), red indicates degradation.
7.2 Performance across Prompting and Planning Methods
@@
Beyond CoT, we also evaluated graph-augmented prompting and interactive planning. These methods show model-dependent effectiveness with no universally superior approach. Full results are provided in Appendix K.